Grade 12th passAlgebra

Fill in the boxes with the correct symbol out of >,
1/-3 [ ] -1/4

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

The LCM of the denominators 3 and 4 is 12
∴(-1/3) = [(-1 × 4)/ (3 × 4)] = (-4/12)
And (-1/4) = [(-1 × 3)/ (4 × 3)] = (-3/12)
Now,
-4 < - 3
So, (-4/12) < (- 3/12)
Hence, 1/-3 [<] -1/4

Thanks